# refrences.bib

### v1

@@ 5.3 Simplification strategies for finding optimal solutions
Based on Theorem 5, as the first simplification strategy, we can accelerate the identification of the optimal solutions set by focusing on the set $\bar{F}$ instead of $F$. In other words, those candidate solutions $x(e)$ (Theorem 4) that belong to $F-\bar{F}$ are not optimal and therefore can be excluded from consideration. However, by using a branch-and-bound method (as the second simplification strategy), we can find an optimal solution without explicitly generating the whole set of solutions $x(e)$ $(e\in\bar{E})$. The branch-and-bound method incrementally builds modified functions $e\in\bar{E}$ by selecting the columns of $\bar{\mathscr{J}}_{i}(e)$ for each row $i$. More precisely, let $n_{i}$ denote the cardinality of $\bar{\mathscr{J}}_{i}(e)$. The branch-and-bound method begins by choosing the columns of $\bar{\mathscr{J}}_{1}(e)$ and generating all partial sequences $e=\left[j_{1}\right]$, $\forall j_{1}\in\bar{\mathscr{J}}_{1}(e)$. Each $e$ is represented by one node, and therefore this yields $n_{1}$ partial branches denoted by nodes 1 to $n_{1}$. Also, for each $e=\left[j_{1}\right]$, we have $\mathscr{I}_{j_{1}}(e)=\{1\}$ and $\mathscr{I}_{j}(e)=\varnothing$, $\forall j\in\mathscr{J}-\left\{j_{1}\right\}$. So, associated with each node $e=\left[j_{1}\right]$, the solution $x(e)$ can be obtained by (15) with the objective value of $cx(e)$. Subsequently, we branch at the node with a smaller objective value. Without loss of generality, suppose that the solution $x\left(e^{*}\right)$, generated by $e^{*}=\left[j_{1}^{*}\right]$, has the smallest objective value. Then, in the second step, we focus on partial sequences $e=\left[j_{1}^{*},j_{2}\right]$, $\forall j_{2}\in\bar{\mathscr{J}}_{2}(e)$, that yields nodes $n_{1}+1$ to $n_{1}+n_{2}$. Similarly, for new nodes $e=\left[j_{1}^{*},j_{2}\right]$ $\left(j_{2}\in\bar{\mathscr{J}}_{2}(e)\right)$, solutions $x(e)$ are attained by (15), and we branch on the new node with a smaller objective value. It should be noted that for successive nodes $e^{\prime}=\left[j_{1},\ldots,j_{p}\right]$ and $e^{\prime\prime}=\left[j_{1},\ldots,j_{p},j_{p+1}\right]$, either $j_{p+1}\in\left\{j_{1},j_{2},\ldots,j_{p}\right\}$ or $j_{p+1}\notin\left\{j_{1},j_{2},\ldots,j_{p}\right\}$. In the former case, we have $\mathscr{I}_{j}\left(e^{\prime}\right)=\mathscr{I}_{j}\left(e^{\prime\prime}\right)$, $\forall j\in\mathscr{J}$, which implies $x\left(e^{\prime}\right)=x\left(e^{\prime\prime}\right)$ (Definition 8). But, in the latter case, $\mathscr{I}_{j_{p+1}}\left(e^{\prime}\right)=\varnothing$, $\mathscr{I}_{j_{p+1}}\left(e^{\prime\prime}\right)=\{p+1\}$ and $\mathscr{I}_{j}\left(e^{\prime}\right)=\mathscr{I}_{j}\left(e^{\prime\prime}\right)$, $\forall j\in\mathscr{J}-\left\{j_{p+1}\right\}$, which together with Definition 8 and Corollary 4 imply $x\left(e^{\prime}\right)_{j_{p+1}}=L_{j_{p+1}}$, $x\left(e^{\prime\prime}\right)_{j_{p+1}}\in\left\{L_{j_{p+1}},U_{j_{p+1}}\right\}$ and $x\left(e^{\prime}\right)_{j}=x\left(e^{\prime\prime}\right)_{j}$, $\forall j\in\mathscr{J}-\left\{j_{p+1}\right\}$; that is, $x\left(e^{\prime}\right)_{j}\leq x\left(e^{\prime\prime}\right)_{j}$, $\forall j\in\mathscr{J}$. So, in any case, we have $cx\left(e^{\prime}\right)\leq cx\left(e^{\prime\prime}\right)$, which means the objective function is non-decreasing on any branches. By the same reasoning as before, we finally attain a node that corresponds to a complete sequence $e_{1}=\left[j_{1}^{*},\ldots,j_{m-1}^{*},j_{m}\right]$ in which $j_{m}\in\bar{\mathscr{J}}_{m}(e)$. Hence, $e_{1}\in\bar{E}$ and then $x\left(e_{1}\right)\in\bar{F}$. The resulting node can be viewed as a candidate solution with $z_{1}=cx\left(e_{1}\right)$. Now, since the objective function is non-decreasing, all nodes whose values cannot be lower than $z_{1}$, can be eliminated from consideration. However, we branch on those nodes so that it is still possible for them to yield a complete sequence $e_{2}$ such that $z_{1}>z_{2}$. Therefore, evaluating some nodes could result in the elimination of many possible solutions or branches from further consideration. Hence, if the solution of the partial branch we are evaluating is larger than the best solution obtained from a complete branch, there is no need to continue evaluating the current branch. The preceding discussion is now summarized as an algorithm.
@@
3Compute sets $I_{ij},S_{ij},I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and each $j\in\mathscr{J}$ (Remark 2 and Table A3).
@@
5       stop; the problem is infeasible (Lemma 4 (a)).
@@
9             Stop; The Problem is Infeasible (Lemma 4 (b)).
@@
12             Apply simplification techniques (Corollaries 7-13) to determine the values of as many variables as possible. Delete the corresponding columns and redundant equations from the problem.
13            Employ the branch-and-bound method on the modified functions $e\in\bar{E}$ to determine the remaining undecided variables.
@@ 6 Numerical example and application
Example 4.
Here we will take a closer look at the problem presented in Examples 1, 2, and 3.
Step 1.
Tables 1 and 2 show all sets $I_{ij}$ and $S_{ij}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively. In both tables, row $i~{}(i\in\mathscr{I})$ corresponds to equation $i$ and column $j~{}(j\in\mathscr{J})$ corresponds to variable $x_{j}$. By Table 1 and Definition 4, all sets $I_{j}$ are shown in Table 3 for each $j\in\mathscr{J}$. Also, all set $S_{ij}^{\prime}$ ($\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J})$ are summarized in Table 4 by using Table 2, Table 3 and Definition 4.
@@
$\{0.4\}$$\{0.4\}$varnothing $\varnothing$ $\{0.6\}$varnothing $\varnothing$ $\{0\}$varnothing $\{1\}$
$\{1\}$varnothing $\varnothing$ $\varnothing$ $\{0,1\}$varnothing $\{1\}$$\{0.2\}$varnothing $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$$\{0.55\}$$\{1\}$varnothing $\varnothing$
$\varnothing$ $\{0.64\}$$\{0.6\}$$\{1\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$$\{1\}$
$\{0.4\}$varnothing $\{0.6\}$$\{1\}$$\{0.3\}$varnothing $\varnothing$ $\{0\}$$\{0\}$$\{0\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.2\}$$\{0.2,0.8\}$varnothing
$\varnothing$ $\{0\}$$\{0,1\}$varnothing $\varnothing$ $\varnothing$ $\{1\}$$\{1\}$varnothing $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,0.8\}$varnothing
$\varnothing$ $\varnothing$ $\{0,1\}$varnothing $\{0\}$varnothing $\varnothing$ $\{0\}$$\{1\}$$\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$varnothing $\{1\}$varnothing $\varnothing$
@@
$\{0.4\}$$\{0.4\}$varnothing $\varnothing$ $\{0.6\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\{1\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.2\}$varnothing $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$$\{0.55\}$$\{1\}$varnothing $\varnothing$
$\varnothing$ $\{0.64\}$$\{0.6\}$$\{1\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\{0.4\}$varnothing $\{0.6\}$$\{1\}$$\{0.3\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.2\}$$\{0.2,0.8\}$varnothing
$\varnothing$ $\varnothing$ $\{0\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\{1\}$varnothing $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.8\}$varnothing
$\varnothing$ $\varnothing$ $\{0\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$varnothing $\{1\}$varnothing $\varnothing$
Step 2.
@@
Step 4.
Since $I_{10}=\{0.6\}$ is a singleton set (see Table 3 ) and $0.6\in S_{7,10}^{\prime}\cap S_{9,10}^{\prime}$ (see Table 4), Corollary 8 indicates that $x_{10}=0.6$ for each feasible solution $x$ (particularly, $x_{10}^{*}=0.6$ for each optimal solution $x^{*}$ ). Also, column 10 and both rows 7 and 9 can be removed from the problem. So, by this simplification technique, the upper bound of the number of admissible functions is reduced from 5184 (Example 2) to $\prod_{i\in\mathscr{I}-\{7,9\}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 3\times 4\times 2\times 1\times 2=864$; that is $|E|\leq 864$.
Moreover, since in Table 4, $S_{10,j}^{\prime}\subseteq S_{3j}^{\prime}$, $\forall j\in\mathscr{J}$, it is concluded that the third equation is a redundant constraint, and it can be deleted by Corollary 9. Hence, by applying this simplification technique, the upper bound of the number of admissible functions is further reduced from 864 to $\prod_{i\in\mathscr{I}-\{3,7,9\}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 4\times 2\times 1\times 2=288$; that is $|E|\leq 288$.
Based on Table 4, it follows that the tenth row itself can also be deleted by Corollary 11; because, $S_{10,6}^{\prime}=\{0,1\}$ and therefore $\left|S_{10,6}^{\prime}\right|=2$. As a result, the above-mentioned upper bound is reduced from 288 to $\prod_{i\in\mathscr{I}-\{3,7,9,10\}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 4\times 2\times 1=144$, and therefore $|E|\leq 144$.
Subsequently, according to Table 4, it turns out that $\mathscr{J}_{8}=\{9\}$ and $S_{89}^{\prime}=\{0.8\}$ are singleton sets and $0.8\in S_{69}^{\prime}$. Thus, $x_{9}^{*}$ is assigned to $x_{9}^{*}=0.8$, and also column 9 and rows 6 and 8 can be deleted by Corollary 10. Consequently, we have $\prod_{i\in\mathscr{I}\{3,6,7,8,10\}}|\mathscr{J}_{i}|=3\times 2\times 3\times 4=72$ and $|E|\leq 72$.
@@
$\{0.4\}$$\{0.4\}$varnothing $\varnothing$ $\{0.6\}$varnothing $\varnothing$ $\varnothing$
$\{1\}$varnothing $\varnothing$ $\varnothing$ $\{1\}$varnothing $\varnothing$ $\{0.2\}$
$\varnothing$ $\{0.64\}$$\{0.6\}$$\{1\}$varnothing $\varnothing$ $\varnothing$ $\varnothing$
$\{0.4\}$varnothing $\{0.6\}$$\{1\}$$\{0.3\}$varnothing $\varnothing$ $\varnothing$
By considering Table 5, it follows that $\mathscr{I}_{8}=\{2\}$ and $S_{28}^{\prime}=\{0.2\}$. So, we have $L_{8}=0.2$ (see Table 3) and therefore $L_{8}\in S_{i8}^{\prime}$, $\forall i\in\mathscr{I}_{8}$. Hence, Corollary 12 implies that $x(e)_{8}=L_{8}=0.2$, $\forall e\in E$. Consequently, by Theorem 4, it is concluded that $x_{8}^{*}=0.2$ for each optimal solution $x^{*}$. Also, column 8 and row 2 can be removed from the problem. So, this simplification technique leads us to $\prod_{i\in\mathscr{I}-\{2,3,6,7,8,9,10\}}\left|\mathscr{J}_{i}\right|=3\times 3\times 4=36$ and $|E|\leq 36$. In addition, since $S_{i6}^{\prime}=S_{i7}^{\prime}=\varnothing$ for each $i\in\{1,2,3,4\}$ (see Table 5), from Definition 5 we have $6\notin\mathscr{J}_{i}$ and $7\notin\mathscr{J}_{i}$, $\forall i\in\{1,2,3,4\}$. Thus, from Definition 6, $\mathscr{I}_{6}(e)=\mathscr{I}_{7}(e)=\varnothing$, $\forall e\in E$, which together with (15) imply $x(e)_{6}=L_{6}$ and $x(e)_{7}=L_{7}$, $\forall e\in E$ (where $L_{6}=L_{7}=0$ from Table 3). So, by using Theorem 4, we can set
$x_{6}^{*}=x_{7}^{*}=0$ and delete columns 6 and 7 . After these reductions, matrices $A^{+}$and $A^{-}$ , vector $b$ and Table 5 are updated as follows:

### v2

@@ 5.3 Simplification strategies for finding optimal solutions
Based on Theorem 5.19, as the first simplification strategy, we can accelerate the identification of the optimal solutions set by focusing on the set $\bar{F}$ instead of $F$. In other words, those candidate solutions $x(e)$ (Theorem 4.2) that belong to $F-\bar{F}$ are not optimal and therefore can be excluded from consideration. However, by using a branch-and-bound method (as the second simplification strategy), we can find an optimal solution without explicitly generating the whole set of solutions $x(e)$ $(e\in\bar{E})$. The branch-and-bound method incrementally builds modified functions $e\in\bar{E}$ by selecting the columns of $\bar{\mathscr{J}}_{i}(e)$ for each row $i$. More precisely, let $n_{i}$ denote the cardinality of $\bar{\mathscr{J}}_{i}(e)$. The branch-and-bound method begins by choosing the columns of $\bar{\mathscr{J}}_{1}(e)$ and generating all partial sequences $e=\left[j_{1}\right]$, $\forall j_{1}\in\bar{\mathscr{J}}_{1}(e)$. Each $e$ is represented by one node, and therefore this yields $n_{1}$ partial branches denoted by nodes 1 to $n_{1}$. Also, for each $e=\left[j_{1}\right]$, we have $\mathscr{I}_{j_{1}}(e)=\{1\}$ and $\mathscr{I}_{j}(e)=\varnothing$, $\forall j\in\mathscr{J}-\left\{j_{1}\right\}$. So, associated with each node $e=\left[j_{1}\right]$, the solution $x(e)$ can be obtained by (12) with the objective value of $cx(e)$. Subsequently, we branch at the node with a smaller objective value. Without loss of generality, suppose that the solution $x\left(e^{*}\right)$, generated by $e^{*}=\left[j_{1}^{*}\right]$, has the smallest objective value. Then, in the second step, we focus on partial sequences $e=\left[j_{1}^{*},j_{2}\right]$, $\forall j_{2}\in\bar{\mathscr{J}}_{2}(e)$, that yields nodes $n_{1}+1$ to $n_{1}+n_{2}$. Similarly, for new nodes $e=\left[j_{1}^{*},j_{2}\right]$ $\left(j_{2}\in\bar{\mathscr{J}}_{2}(e)\right)$, solutions $x(e)$ are attained by (12), and we branch on the new node with a smaller objective value. It should be noted that for successive nodes $e^{\prime}=\left[j_{1},\ldots,j_{p}\right]$ and $e^{\prime\prime}=\left[j_{1},\ldots,j_{p},j_{p+1}\right]$, either $j_{p+1}\in\left\{j_{1},j_{2},\ldots,j_{p}\right\}$ or $j_{p+1}\notin\left\{j_{1},j_{2},\ldots,j_{p}\right\}$. In the former case, we have $\mathscr{I}_{j}\left(e^{\prime}\right)=\mathscr{I}_{j}\left(e^{\prime\prime}\right)$, $\forall j\in\mathscr{J}$, which implies $x\left(e^{\prime}\right)=x\left(e^{\prime\prime}\right)$ (Definition 4.1). But, in the latter case, $\mathscr{I}_{j_{p+1}}\left(e^{\prime}\right)=\varnothing$, $\mathscr{I}_{j_{p+1}}\left(e^{\prime\prime}\right)=\{p+1\}$ and $\mathscr{I}_{j}\left(e^{\prime}\right)=\mathscr{I}_{j}\left(e^{\prime\prime}\right)$, $\forall j\in\mathscr{J}-\left\{j_{p+1}\right\}$, which together with Definition 4.1 and Corollary 2.16 imply $x\left(e^{\prime}\right)_{j_{p+1}}=L_{j_{p+1}}$, $x\left(e^{\prime\prime}\right)_{j_{p+1}}\in\left\{L_{j_{p+1}},U_{j_{p+1}}\right\}$ and $x\left(e^{\prime}\right)_{j}=x\left(e^{\prime\prime}\right)_{j}$, $\forall j\in\mathscr{J}-\left\{j_{p+1}\right\}$; that is, $x\left(e^{\prime}\right)_{j}\leq x\left(e^{\prime\prime}\right)_{j}$, $\forall j\in\mathscr{J}$. So, in any case, we have $cx\left(e^{\prime}\right)\leq cx\left(e^{\prime\prime}\right)$, which means the objective function is non-decreasing on any branches. By the same reasoning as before, we finally attain a node that corresponds to a complete sequence $e_{1}=\left[j_{1}^{*},\ldots,j_{m-1}^{*},j_{m}\right]$ in which $j_{m}\in\bar{\mathscr{J}}_{m}(e)$. Hence, $e_{1}\in\bar{E}$ and then $x\left(e_{1}\right)\in\bar{F}$. The resulting node can be viewed as a candidate solution with $z_{1}=cx\left(e_{1}\right)$. Now, since the objective function is non-decreasing, all nodes whose values cannot be lower than $z_{1}$, can be eliminated from consideration. However, we branch on those nodes so that it is still possible for them to yield a complete sequence $e_{2}$ such that $z_{1}>z_{2}$. Therefore, evaluating some nodes could result in the elimination of many possible solutions or branches from further consideration. Hence, if the solution of the partial branch we are evaluating is larger than the best solution obtained from a complete branch, there is no need to continue evaluating the current branch. The preceding discussion is now summarized as an algorithm.
@@
3Compute sets $I_{ij},S_{ij},I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and each $j\in\mathscr{J}$ (Remark 2.13 and Table A3).
@@
5   stop; the problem is infeasible (Lemma 3.1 (a)).
@@
9      Stop; The Problem is Infeasible (Lemma 3.1 (b)).
@@
12      Apply simplification techniques (Corollaries 5.3-5.16).
13     Delete the corresponding columns and redundant equations from the problem.
14     Employ the branch-and-bound method on the modified functions $e\in\bar{E}$.
@@
In the first step of the algorithm, the $I_{ij},S_{ij},I_{j}$ and $S_{ij}^{\prime}$ sets are computed using at most $O(mn)$ comparisons with the aid of Remark 2.13 and Table A3. In addition, the conditions in steps 2 and 6 of algorithm involve a comparison of at most $O(mn)$. It may be necessary to perform up to $O(m(m+1)/2)$ operations in order to apply the simplification techniques in step 10 of algorithm. There is, however, a possibility that the simplification process may exhibit non-polynomial complexity if the problem variables exceed a polynomial growth rate. Similarly, step 11 of algorithm requires up to $O(mn)$ operations (considering that $m$ and $n$ have been modified as a result of the simplification techniques). In step 12 of algorithm, if the upper bound of the admissible function $|E|$ exceeds polynomial complexity, as mentioned in the introduction, the problem falls into the NP-hard complexity class of computation. In addition, Flowchart 1 provides a better understanding of Algorithm 1.
According to the algorithm described in this paper, we are able to obtain an exact (non-numerical and non-probabilistic) solution to Problem 1. Unlike other studies that are focused on nonlinear objective functions, this paper examines a linear problem with constraints defined by bipolar fuzzy equations using continuous Archimedean t-norms. It is therefore not possible to analyze algorithms using benchmarks in the same manner as numerical or heuristic methods. Moreover, the study proved theorems, lemmas, and results common to all continuous Archimedean t-norms relevant to Problem 1 (Table A1 contains a list of common continuous Archimedean t-norms). In consequence, variation in the t-norms does not significantly alter the core steps of Algorithm 1; rather, the differences lie in their properties and characteristics. According to the findings of the current study, future work may focus on enhancing the proposed algorithm or further simplifying Problem 1, which would collectively contribute to the improved practical performance of the algorithm.
Figure 1: Flowchart of Algorithm 1
@@ 6 Numerical example and application
Example 6.1.
Here we will take a closer look at the problem presented in Examples 2.17, 3.11, and 5.18.
Step 6.2.
Tables 1 and 2 show all sets $I_{ij}$ and $S_{ij}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively. In both tables, row $i~(i\in\mathscr{I})$ corresponds to equation $i$ and column $j~(j\in\mathscr{J})$ corresponds to variable $x_{j}$. By Table 1 and Definition 2.15, all sets $I_{j}$ are shown in Table 3 for each $j\in\mathscr{J}$. Also, all set $S_{ij}^{\prime}$ ($\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J})$ are summarized in Table 4 by using Table 2, Table 3 and Definition 2.15.
@@
$\{0.4\}$ $\{0.4\}$ $\varnothing$ $\varnothing$ $\{0.6\}$ $\varnothing$ $\varnothing$ $\{0\}$ $\varnothing$ $\{1\}$
$\{1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$ $\varnothing$ $\{1\}$ $\{0.2\}$ $\varnothing$ $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$ $\{0.55\}$ $\{1\}$ $\varnothing$ $\varnothing$
$\varnothing$ $\{0.64\}$ $\{0.6\}$ $\{1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$ $\{1\}$
$\{0.4\}$ $\varnothing$ $\{0.6\}$ $\{1\}$ $\{0.3\}$ $\varnothing$ $\varnothing$ $\{0\}$ $\{0\}$ $\{0\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.2\}$ $\{0.2,0.8\}$ $\varnothing$
$\varnothing$ $\{0\}$ $\{0,1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\{1\}$ $\{1\}$ $\varnothing$ $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,0.8\}$ $\varnothing$
$\varnothing$ $\varnothing$ $\{0,1\}$ $\varnothing$ $\{0\}$ $\varnothing$ $\varnothing$ $\{0\}$ $\{1\}$ $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$ $\varnothing$ $\{1\}$ $\varnothing$ $\varnothing$
@@
$\{0.4\}$ $\{0.4\}$ $\varnothing$ $\varnothing$ $\{0.6\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\{1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.2\}$ $\varnothing$ $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$ $\{0.55\}$ $\{1\}$ $\varnothing$ $\varnothing$
$\varnothing$ $\{0.64\}$ $\{0.6\}$ $\{1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\{0.4\}$ $\varnothing$ $\{0.6\}$ $\{1\}$ $\{0.3\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.2\}$ $\{0.2,0.8\}$ $\varnothing$
$\varnothing$ $\varnothing$ $\{0\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{1\}$ $\varnothing$ $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.8\}$ $\varnothing$
$\varnothing$ $\varnothing$ $\{0\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0.6\}$
$\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ $\{0,1\}$ $\varnothing$ $\{1\}$ $\varnothing$ $\varnothing$
Step 6.3.
@@
Step 6.5.
Since $I_{10}=\{0.6\}$ is a singleton set (see Table 3) and $0.6\in S_{7,10}^{\prime}\cap S_{9,10}^{\prime}$ (see Table 4), Corollary 5.5 indicates that $x_{10}=0.6$ for each feasible solution $x$ (particularly, $x_{10}^{*}=0.6$ for each optimal solution $x^{*}$ ). Also, column 10 and both rows 7 and 9 can be removed from the problem. So, by this simplification technique, the upper bound of the number of admissible functions is reduced from 5184 (Example 3.11) to $\prod_{i\in\mathscr{I}-\{7,9\}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 3\times 4\times 2\times 1\times 2=864$; that is $|E|\leq 864$.
Moreover, since in Table 4, $S_{10,j}^{\prime}\subseteq S_{3j}^{\prime}$, $\forall j\in\mathscr{J}$, it is concluded that the third equation is a redundant constraint, and it can be deleted by Corollary 5.7. Hence, by applying this simplification technique, the upper bound of the number of admissible functions is further reduced from 864 to $\prod_{i\in\mathscr{I}-\{3,7,9\}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 4\times 2\times 1\times 2=288$; that is $|E|\leq 288$.
Based on Table 4, it follows that the tenth row itself can also be deleted by Corollary 5.11; because, $S_{10,6}^{\prime}=\{0,1\}$ and therefore $\left|S_{10,6}^{\prime}\right|=2$. As a result, the above-mentioned upper bound is reduced from 288 to $\prod_{i\in\mathscr{I}-\{3,7,9,10\}}\left|\mathscr{J}_{i}\right|=3\times 2\times 3\times 4\times 2\times 1=144$, and therefore $|E|\leq 144$.
Subsequently, according to Table 4, it turns out that $\mathscr{J}_{8}=\{9\}$ and $S_{89}^{\prime}=\{0.8\}$ are singleton sets and $0.8\in S_{69}^{\prime}$. Thus, $x_{9}^{*}$ is assigned to $x_{9}^{*}=0.8$, and also column 9 and rows 6 and 8 can be deleted by Corollary 5.9. Consequently, we have $\prod_{i\in\mathscr{I}-\{3,6,7,8,9,10\}}|\mathscr{J}_{i}|=3\times 2\times 3\times 4=72$ and $|E|\leq 72$.
@@
$\{0.4\}$ $\{0.4\}$ $\varnothing$ $\varnothing$ $\{0.6\}$ $\varnothing$ $\varnothing$ $\varnothing$
$\{1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\{1\}$ $\varnothing$ $\varnothing$ $\{0.2\}$
$\varnothing$ $\{0.64\}$ $\{0.6\}$ $\{1\}$ $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$
$\{0.4\}$ $\varnothing$ $\{0.6\}$ $\{1\}$ $\{0.3\}$ $\varnothing$ $\varnothing$ $\varnothing$
By considering Table 5, it follows that $\mathscr{I}_{8}=\{0.2\}$ and $S_{28}^{\prime}=\{0.2\}$. So, we have $L_{8}=0.2$ (see Table 3) and therefore $L_{8}\in S_{i8}^{\prime}$, $\forall i\in\mathscr{I}_{8}$. Hence, Corollary 5.14 implies that $x(e)_{8}=L_{8}=0.2$, $\forall e\in E$. Consequently, by Theorem 4.2, it is concluded that $x_{8}^{*}=0.2$ for each optimal solution $x^{*}$. Also, column 8 and row 2 can be removed from the problem. So, this simplification technique leads us to $\prod_{i\in\mathscr{I}-\{2,3,6,7,8,9,10\}}\left|\mathscr{J}_{i}\right|=3\times 3\times 4=36$ and $|E|\leq 36$. In addition, since $S_{i6}^{\prime}=S_{i7}^{\prime}=\varnothing$ for each $i\in\{1,2,3,4\}$ (see Table 5), from Definition 3.5 we have $6\notin\mathscr{J}_{i}$ and $7\notin\mathscr{J}_{i}$, $\forall i\in\{1,2,3,4\}$. Thus, from Definition 3.6, $\mathscr{I}_{6}(e)=\mathscr{I}_{7}(e)=\varnothing$, $\forall e\in E$, which together with (12) imply $x(e)_{6}=L_{6}$ and $x(e)_{7}=L_{7}$, $\forall e\in E$ (where $L_{6}=L_{7}=0$ from Table 3). So, by using Theorem 4.2, we can set $x_{6}^{*}=x_{7}^{*}=0$ and delete columns 6 and 7 . After these reductions, matrices $A^{+}$and $A^{-}$, vector $b$ and Table 5 are updated as follows: